OpenStudy (anonymous):

explain how to solve this?: Five years ago, a man was seven times as old as his niece was at that time. 10 yrs from now, the man will be twice as old as his niece will then be. Find the present ages of the man and of the girl.

OpenStudy (anonymous):

oh god, please. is it that hard? D:

OpenStudy (anonymous):

fff. i really need this dear god. i just need an idea of what it is they want me to do. they didn't give me any numbers (i'm sure that's the point, but where am i supposed to get the numbers from. grrrrr. what is it, like 7x-5 2x=10? D: D:

OpenStudy (anonymous):

this must be an advanced question. I understand all of the other questions on here. maybe i'm just on the wrong sight. >.<

OpenStudy (anonymous):

Sorry for taking so long! It's been a while since I've done problems like these. Let X = the man's current age; Y = the girl's age back then X-5 = 7Y X+10 = 2(Y+15) Solve for X in terms of Y or Y in terms of X. Man's present age = 26, girl's present age = 8 Sorry again!